Keshawn works at a party-planning company and makes $7 per hour. His weekly paycheck also includes a $12 bonus for every party he schedules. His $176 paycheck last week was for 20 hours. How many parties did he schedule? Explain how you found your answer
Answer by mananth(16949)

Keshawn works at a party-planning company and makes $7 per hour. His weekly paycheck also includes a $12 bonus for every party he schedules. His $176 paycheck last week was for 20 hours. How many parties did he schedule? Explain how you found your answer
...
number of parties be x
bonus on parties = 12x
..
rate $7 /hour
20 hours
20*7 =$140 for hours put in.
140+12x=176
-140 -140
12x=36
/12
x=3 parties in the week.
...
Check
..
20*7 +3*12 = 176
..
m.ananth@hotmail.ca